# Multiplication of 0-1 matrices via clustering

### v2

@@ 5 Extensions
The rows or columns in the input 0-1 matrices can be very long. Also, a large number of clusters might be needed in order to obtain a low upper bound on their radius. Among other things, for these reasons, we have picked Gonzalez’s classical algorithm for the $k$-center clustering problem [16] as a basic tool in our approach to the arithmetic matrix product of two 0-1 matrices with clustered rows or columns. The running time of his algorithm is linear not only in the number of input points but also in their dimension, and in the parameter $k.$ Importantly, it is very simple and provides a solution within $2$ of the optimum. For instance, there exist faster (in terms of $n$ and $k$) $2$-approximation algorithms for $k$-center clustering with hidden exponential dependence on the dimension in their running time, see [10, 17],
@@
One could easily generalize our main results by replacing Gonzalez’s algorithm with a simple and efficient approximation algorithm for the more general problem of $k$-center clustering with outliers [6]. In the latter problem, a given number $z$ of input points could be discarded as outliers when trying to minimize the maximum cluster radius. Unfortunately, the algorithms for this more general problem tend to be more complicated and the focus seems to be the approximation ratio achievable in polynomial time (e.g., $3$ in [6] and $2$ in [18]) not the time complexity.
@@
There are many other variants of clustering than $k$-center clustering, and plenty of methods have been developed for them in the literature. In fact, in the design of efficient algorithms for the exact arithmetic matrix product of 0-1 matrices with clustered rows or columns, using the $k$-median clustering could seem more natural. The objective in the latter problem is to minimize the sum of distances between the input points and their nearest centers. Unfortunately, no simple deterministic $O(1)$-approximation algorithms for the latter problem that are efficient in case the dimension and $k$ parameters are large seem be to available [5, 7].
@@
Acknowledgments
J.J. was partially supported by KAKENHI grant 24K22294.

### v3

@@ 5 Potential extensions
The rows or columns in the input 0-1 matrices can be very long. Also, a large number of clusters might be needed in order to obtain a low upper bound on their radius. Among other things, for these reasons, we have picked Gonzalez’s classical algorithm for the $k$-center clustering problem [Gon85] as a basic tool in our deterministic approach to the arithmetic matrix product of two 0-1 matrices with clustered rows or columns. The running time of his algorithm is linear not only in the number of input points but also in their dimension, and in the parameter $k.$ Importantly, it is very simple, deterministic, provides a solution within $2$ of the optimum, and can be applied in Hamming spaces. For instance, there exist faster (in terms of $n$ and $k$) 22-approximation algorithms for $k$-center clustering with hidden exponential dependence on the dimension in their running time, see [FG88, HM06]. Among several newer works on speeding approximation algorithms for $k$-center clustering (e.g., [EHS20, FJLNP25, JKS24, KLP26]) only [KLP26] explicitly includes Hamming spaces. The randomized approximation method for the $k$-center clustering problem in Hamming spaces from [KLP26] is summarized in Fact 2. It provides approximation guarantees arbitrarily close to $2,$ is substantially faster than Gonzalez’s algorithm when the dimension and $k$ are superlogarithmic, and is relatively simple. For these reasons, our improved randomized algorithms for the approximate and exact 0-1 matrix multiplication are based on this method.
@@
One could easily generalize our main results by replacing approximation algorithms for $k$-center clustering with approximation algorithms for the more general problem of $k$-center clustering with outliers [Char01]. In the latter problem, a given number $z$ of input points could be discarded as outliers when trying to minimize the maximum cluster radius. Unfortunately, the algorithms for this more general problem tend to be more complicated and the focus seems to be on the approximation ratio achievable in polynomial time (e.g., $3$ in [Char01] and $2$ in [lot]) and not on the time complexity.
@@
There are many other variants of clustering than $k$-center clustering, and plenty of methods have been developed for them in the literature. In fact, in the design of efficient algorithms for the exact arithmetic matrix product of 0-1 matrices with clustered rows or columns, using the $k$-median clustering could seem more natural. The objective in the latter problem is to minimize the sum of distances between the input points and their nearest centers. Unfortunately, no simple deterministic $O(1)$-approximation algorithms for the latter problem that are efficient in case the dimension and $k$ parameters are large seem be to available [char99, chen].
@@
Funding acknowledgments
Jesper Jansson was partially supported by Japan Society for the Promotion of Science - KAKENHI grant 24K22294.